# Near Coincidences and Nilpotent Division Fields

### v1

@@ 4.3. Modular curves associated to split Cartan subgroups
The work of Bilu, Parent, and Rebolledo work in [3] gives an almost complete picture of rational points on $X_{s}^{+}(p)$. This work together with the work of Balakrishnan, Dogra, and Müller in [1] gives, among other things, the following theorem.
@@
[1, 3] If $p\geq 11$ is a prime, then the rational points on $X_{s}^{+}(p)$ are cusps or correspond to elliptic curves with complex multiplication.
The only cases that remain would be to classify when $E/\mathbb{Q}$ can have its mod $p$ image contained in the normalizer of a split Cartan for $p=2$, $3$, $5$, and $7$. For these values of $p$, $X_{s}^{+}(p)$ is isomorphic to $\mathbb{P}^{1}$ (see [26, 33, 36] among other places).
The case when $p=2$ is a simple one since all of the proper subgroups of $\operatorname{GL}_{2}(\mathbb{Z}/2\mathbb{Z}){\simeq S_{3}}$ are abelian, while $S_{3}$ is not nilpotent. When $p$ is a Fermat prime, $|C_{s}^{+}(p)|=2(p-1)^{2}$ has order a power of $2$, and so in these case the image of $\bar{\rho}_{E,p}$ is nilpotent. When $p=7$, the projective image of $C_{s}^{+}(p)$ has size $2(7-1)=12$ and so is not nilpotent. As shown in [26], there are three maximal admissible subgroups of $C_{s}^{+}(7)$, and for each of these, the corresponding modular curve has genus $1$. For two of these, there are no non-CM points, while for the third of these, there is a non-CM point with $j=\frac{3^{3}\cdot 5\cdot 7^{5}}{2^{7}}$. An elliptic curve with this $j$-invariant has mod $7$ image isomorphic to either $\mathbb{Z}/6\mathbb{Z}\times S_{3}$ or $\mathbb{Z}/3\mathbb{Z}\times S_{3}$ and neither of these groups is nilpotent.
This subsection can be summarized in the following proposition.
@@
Let $E/\mathbb{Q}$ be an elliptic curve without complex multiplication and let $p$ be an odd prime such that $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in $C_{s}^{+}(p)$. If $\mathbb{Q}(E[p])/\mathbb{Q}$ is a nilpotent extension, then $p=3$ or $5$.
@@ 4.4. Modular curves associated to non-split Cartan subgroups
where $\epsilon$ is a generator of $(\mathbb{Z}/n\mathbb{Z})^{\times}$. The normalizer of this group in $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is exactly
$C_{ns}^{+}(p):=\left\langle C_{ns}^{+}(p),\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\right\rangle.$
We will denote the modular curves corresponding to these groups by $X_{ns}(p)$ and $X_{ns}^{+}(p)$.
Much less is known about the rational points on the modular curves associated with the normalizers of the non-split Cartan subgroups. These modular curves have some arithmetic properties that make analysis of their rational points particularly challenging. In particular, the Jacobians of these modular curves always have rank at least as big as the genus of the curve. This rules out the traditional method of Chabauty and Coleman and requires more advanced techniques (which have been successful in two cases: see [1] and [2]). Fortunately for us, enough is known that we will be able to say quite a bit about the situation unconditionally, and the remainder of what we need is covered by Conjecture 1.6.
@@
[36, Proposition 1.13] Let $E/\mathbb{Q}$ be an elliptic curve that does not have complex multiplication and let $p\geq 17$ be a prime such that $\bar{\rho}_{E,p}$ is not surjective and for which $E$ does not have a $\mathbb{Q}$-rational cyclic $p$-isogeny.
  • If $p\equiv 1\bmod 3$, then $\operatorname{Im}\bar{\rho}_{E,p}$ is conjugate to $C_{ns}^{+}(p)$ in $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$.
  • If $p\equiv 2\bmod 3$, then $\operatorname{Im}\bar{\rho}_{E,p}$ is conjugate to with $C_{ns}^{+}(p)$ or
    $G(p):=\{a^{3}\colon a\in C_{ns}(p)\}\cup\{a^{3}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\colon a\in C_{ns}(p)\}$
    in $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z}).$
Theorem 4.13.
[13, Theorem 1.6] The only prime $p\neq 2,$ or $3$ such that there is an elliptic curve $E/\mathbb{Q}$ without complex multiplication such that $\operatorname{Im}\bar{\rho}_{E,p}$ is conjugate to $G(p)$ is $p=5.$
From this we can see that if $E/\mathbb{Q}$ is an elliptic curve and $p\geq 7$ is a prime such that $\operatorname{Im}\bar{\rho}_{E,p}$ is conjugate to a subgroup of $C_{ns}^{+}(p),$ then $\operatorname{Im}\bar{\rho}_{E,p}=C_{ns}^{+}(p)$ and the image of $\bar{\rho}_{E,p}$ in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ has size
@@
Combining this with Example 4.6, we get that in this case $\operatorname{Im}\bar{\rho}_{E,p}$ is nilpotent exactly when $2(p+1)$ is a power of 2 which can happen only when $p+1$ is a power of 2 or $p$ is a Mersenne prime.
@@
Let $E/\mathbb{Q}$ be an elliptic curve and let $p$ be a prime such that $\operatorname{Im}\bar{\rho}_{E,p}$ is conjugate to a subgroup of $C_{ns}^{+}(p)$ and $\mathbb{Q}(E[p])/\mathbb{Q}$ is a nilpotent extension. Then $p$ is a Mersenne prime.
@@ 4.5. The case of complex multiplication
One of the interesting properties of elliptic curves with complex multiplication is that their mod $p$ representations almost always have their images in a Cartan subgroup. From the discussion in [5, Pages 194-195] we get the following theorem.
Theorem 4.17.
Let $E/\mathbb{Q}$ be an elliptic curve with complex multiplication by $\mathcal{O}$ an order of $K$ and let $p\geq 5$ be a prime such that $p\nmid{\rm disc}(\mathcal{O})$. Let $\mathcal{O}_{K}$ be the ring of integers of $K.$ If the ideal $p\mathcal{O}_{K}$ splits in $\mathcal{O}_{K}$, then $\operatorname{Im}\bar{\rho}_{E,p}$ is conjugate to a subgroup of $C_{s}^{+}(p)$. On the other hand, if the ideal $p\mathcal{O}_{K}$ is inert in $\mathcal{O}_{K}$, then $\operatorname{Im}\bar{\rho}_{E,p}$ is conjugate to a subgroup of $C_{ns}^{+}(p)$.
In fact, we can say a little more than what Theorem 4.17 says. Proposition 1.14(i) and (ii) of [36] states the following.
@@
$\begin{cases}C_{s}^{+}(p)&\hbox{if $p\mathcal{O}_{K}$ splits in $\mathcal{O}_{K}$, and}\\ C_{ns}^{+}(p)&\hbox{if $p\mathcal{O}_{K}$ is inert in $\mathcal{O}_{K}$}.\end{cases}$
The point of this proposition is that in these cases the mod $p$ images is as large as possible. This is useful because we know that in these cases $\operatorname{Im}\bar{\rho}_{E,p}$ is its image in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is genuinely dihedral and has easily computable size. From all the work we have done up till now, we know that that $\operatorname{Im}\bar{\rho}_{E,p}$ will be nilpotent exactly when its image in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is a 2-group.
@@
The discussion proceeding this theorem covers the case where $p\nmid{\rm disc}(\mathcal{O}).$ To handle the case where $p\mid{\rm disc}(\mathcal{O}),$ we refer to [36, Theorem 1.14] which shows that in this case $\operatorname{Im}\bar{\rho}_{E,p}$ isomorphic to one of the following groups
@@
In all 3 cases, the matrix $\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$ is in $\operatorname{Im}\bar{\rho}_{E,p}$ and so $\operatorname{Im}\bar{\rho}_{E,p}$ is not nilpotent by the discussion after Remark 4.8. ∎
@@ 4.5.1. The case when $E$ has complex multiplication by $\mathbb{Z}[\zeta_{3}]$
The images of the mod $p$ representations of $E$ depend on the value of $d$ modulo 6th powers. This relationship is explicitly classified in [36, Propositions 1.15 and 1.16]. We summerize the relevant parts of those propositions here for the convenience of the reader.
@@
[36, Propositions 1.15 and 1.16] Let $E/\mathbb{Q}$ be an elliptic curve with complex multiplication by $\mathbb{Z}[\zeta_{3}]$. Then the curve $E$ can be given by a Weierstrass equation of the form
@@
    If $d$ is a cube, then $\operatorname{Im}\bar{\rho}_{E,2}=\left\langle\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right\rangle$. Otherwise, $\operatorname{Im}\bar{\rho}_{E,2}=\operatorname{GL}_{2}(\mathbb{Z}/2\mathbb{Z})$
@@
    If $-4d$ is not a cube, then $\operatorname{Im}\bar{\rho}_{E,3}$ is conjugate to
    {{(±$\begin{cases}\left\{\begin{pmatrix}\pm 1&a\\ 0&b\end{pmatrix}\colon a\in\mathbb{Z}/3\mathbb{Z}\hbox{ and }b\in(\mathbb{Z}/3\mathbb{Z})^{\times}\right\}&\text{ if neither }d\text{ nor }-3d\text{ is a square }\\ \left\{\begin{pmatrix}1&a\\ 0&b\end{pmatrix}\colon a\in(\mathbb{Z}/3\mathbb{Z})^{\times}\hbox{ and }b\in\mathbb{Z}/3\mathbb{Z}\right\}&\text{ if }d\text{ is a square }\\ \left\{\begin{pmatrix}a&b\\ 0&1\end{pmatrix}\colon a\in(\mathbb{Z}/3\mathbb{Z})^{\times}\hbox{ and }b\in\mathbb{Z}/3\mathbb{Z}\right\}&\text{ if }3d\text{ is a square. }\end{cases}$
    On the other hand, if $-4d$ is a cube, then $\operatorname{Im}\bar{\rho}_{E,3}$ is conjugate to
    $\begin{cases}\left\{\begin{pmatrix}a&0\\ 0&b\end{pmatrix}\colon a,b\in(\mathbb{Z}/3\mathbb{Z})^{\times}\right\}&\text{ if neither }d\text{ nor }-3d\text{ is a square }\\ \left\{\begin{pmatrix}1&0\\ 0&b\end{pmatrix}\colon b\in(\mathbb{Z}/3\mathbb{Z})^{\times}\right\}&\text{ ifeither }d\text{ or }-3d\text{ is a square. }\end{cases}$
@@
    $\begin{pmatrix}a&0\\ 0&b\end{pmatrix}\hbox{ and }\begin{pmatrix}0&a\\ b&0\end{pmatrix}$
@@
    Suppose that $p\equiv 2$ or $5\bmod 9$ and let $e\in\{1,2\}$ such that $-e\equiv\frac{p+1}{3}\bmod 3$. If $d\not\equiv 16p^{e}\bmod(\mathbb{Q}^{\times})^{3}$, then $\operatorname{Im}\bar{\rho}_{E,p}$ is conjugate to $C_{ns}^{+}(p)$. If $d\equiv 16p^{e}\bmod(\mathbb{Q}^{\times})^{3}$, then $\operatorname{Im}\bar{\rho}_{E,p}$ is conjugate in $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ to the unique index 3 subgroup of $C_{ns}^{+}(p)$ generated by $C_{ns}(p)$ and $\begin{pmatrix}1&0\\ 0&1\end{pmatrix}.$
The take away from this theorem is that the images of the Galois representations associated to the elliptic curve $E_{d}\colon y^{2}=x^{3}+d$ is completely controlled by $d\bmod(\mathbb{Q}^{\times})^{3}.$
For example, condition (1) tells us that $\mathbb{Q}(E_{d}[2])/\mathbb{Q}$ is a nilpotent extension exactly when $d$ is a cube. Similarly, condition (2) says that $\mathbb{Q}(E_{d}[3])/\mathbb{Q}$ is nilpotent when $-4d$ is a cube.
We note that there are no Fermat primes $\equiv 1\pmod{9}$ and so condition (3) never yields a nilpotent $\mathbb{Q}(E[p])/\mathbb{Q}$. Likewise, there are no Mersenne primes $p\equiv 8\pmod{9}$.
The last cases that we have to deal with are the special cases that arise in cases (5) and (6) of Theorem 4.20. In cases (5) and (6) repsectively, the image of $\bar{\rho}_{E,p}$ is contained in an index 3 subgroup of $C_{s}^{+}(p)$ and $C_{ns}^{+}(p)$ respectively. If we are in condition (5), then image of $\operatorname{Im}\bar{\rho}_{E,p}$ inside of $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is a dihedral group of size $\frac{2(p-1)}{3}$, while in condition (6) the image of $\operatorname{Im}\bar{\rho}_{E,p}$ inside of $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is a dihedral group of size $\frac{2(p+1)}{3}.$ This along with our previous analysis give the following proposition.
@@
$\begin{cases}d\equiv 1\pmod{(\mathbb{Q}^{\times})^{3}}\text{ if }p=2,\\ d\equiv 2\pmod{(\mathbb{Q}^{\times})^{3}}\text{ if }p=3,\\ d\equiv 2\cdot p^{\frac{p-1}{3}}\pmod{(\mathbb{Q}^{\times})^{3}}\text{ if }p=3\cdot 2^{k}+1\hbox{ for some $k\geq 1$},\\ d\equiv 2\cdot p^{\frac{p+1}{3}}\pmod{(\mathbb{Q}^{\times})^{3}}\text{ if }p=3\cdot 2^{k}-1\hbox{ for some $k\geq 1$}.\\ \end{cases}$
@@
The proof of Corollary 1.8 follows from a summary observation. Namely, we observe from this section that if $E/\mathbb{Q}$ is an elliptic curve and $p$ is an odd prime such that $\mathbb{Q}(E[p])/\mathbb{Q}$ is a nilpotent extension, then either $p$ must be of the form $2^{k}\pm 1$ or $3\cdot 2^{k}\pm 1$. The prime $p=19$ is the first prime not in any of these forms.
In Table 1 we list the modular curves that we need to consider and determine which of them have rational points.
$p$ RSZB Label Common Label Has Rational Points
2 2.2.0.1 $X_{ns}(2)$ Yes
2.3.0.1 $X_{0}^{+}(2)$ Yes
3 3.3.0.1 $X_{ns}^{+}(3)$ Yes
5 5.15.0.1 $X_{s}^{+}(5)$ Yes
5.20.0.2 $X_{ns}^{+}(5)$ No
7 7.21.0.1 $X_{ns}^{+}(7)$ Yes
7.56.1.1 $X_{s}^{+}(7)$ No
Table 1. Maximal admissible subgroups of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ to be considered assuming Conjecture 1.6.

### v3

@@ 4.3. Modular curves associated to split Cartan subgroups
The work of Bilu, Parent, and Rebolledo in [3] gives an almost complete picture of rational points on $X_{s}^{+}(p)$. This work together with the work of Balakrishnan, Dogra, Müller, Tuitman and Vonk in [1] gives, among other things, the following theorem.
@@
[1, 3] If $p\geq 11$ is a prime, then the $\mathbb{Q}$-rational points on $X_{s}^{+}(p)$ are cusps or correspond to elliptic curves with complex multiplication.
Using this result we prove the following.
@@
Let $E/\mathbb{Q}$ be an elliptic curve without complex multiplication and let $p$ be a prime such that $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in $C_{s}^{+}(p)$. If $\mathbb{Q}(E[p])/\mathbb{Q}$ is a nilpotent extension, then $p\in\{2,3,5\}$.
Proof.
By Theorem 4.11, it suffices to rule out the case that $p=7$. When $p=7$, the projective image of $C_{s}^{+}(p)$ has size 2(71)=122(7-1)=12 and so is not nilpotent. As shown in [33], there are three maximal admissible subgroups of $C_{s}^{+}(7)$, and for each of these, the corresponding modular curve has genus $1$. For two of these, there are no non-CM points, while for the third of these, there is a non-CM point with $j=\frac{3^{3}\cdot 5\cdot 7^{5}}{2^{7}}$. An elliptic curve with this $j$-invariant has mod $7$ image isomorphic to either $\mathbb{Z}/6\mathbb{Z}\times S_{3}$ or $\mathbb{Z}/3\mathbb{Z}\times S_{3}$ and neither of these groups is nilpotent. ∎
@@ 4.4. Modular curves associated to non-split Cartan subgroups
where $\epsilon$ is a generator of $(\mathbb{Z}/p\mathbb{Z})^{\times}$. The normalizer of this group in $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is exactly
$C_{ns}^{+}(p):=\left\langle C_{ns}(p),\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\right\rangle$.
We will denote the modular curves corresponding to these groups by $X_{ns}(p)$ and $X_{ns}^{+}(p)$respectively.
Much less is known about the rational points on the modular curves associated with the normalizers of the non-split Cartan subgroups compared to what is know about the split Cartan cases. These modular curves have some arithmetic properties that make analysis of their rational points particularly challenging. In particular, the Jacobians of these modular curves always have analytic rank at least as big as the genus of the curve. This rules out the traditional method of Chabauty and Coleman and requires more advanced techniques (which have been successful in two cases: see [1] and [2]). Fortunately for us, enough is known that we will be able to say quite a bit about the situation unconditionally, and the remainder of what we need is covered by Conjecture 1.8.
@@
Let $E/\mathbb{Q}$ be an elliptic curve that does not have complex multiplication and let $p\geq 7$ be a prime such that $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in $C_{ns}^{+}(p)$. Then $\operatorname{Im}\bar{\rho}_{E,p}=C_{ns}^{+}(p)$.
Proof.
In [45] the stated result is proven for $p=7$ (Theorem 1.5) and $p=11$ (Theorem 1.6). For $p=13$, [1] shows that there are no non-CM elliptic curves for which $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in $C_{ns}^{+}(p)$. The cases that $p\geq 17$ are handled by combining Proposition 1.13 of [45] with Theorem 1.6 of [14]. ∎
From this we can see that if $E/\mathbb{Q}$ is an elliptic curve and $p\geq 7$ is a prime such that $\operatorname{Im}\bar{\rho}_{E,p}$ is conjugate to a subgroup of $C_{ns}^{+}(p),$ then $\operatorname{Im}\bar{\rho}_{E,p}=C_{ns}^{+}(p)$ and the image of $\bar{\rho}_{E,p}$ in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is a dihedral group with order
@@
Combining this with Example 4.6 and Remark 4.9, we get that in this case $\operatorname{Im}\bar{\rho}_{E,p}$ is nilpotent exactly when $2(p+1)$ is a power of 2 which can happen only when $p+1$ is a power of 2 or $p$ is a Mersenne prime.
@@
Let $E/\mathbb{Q}$ be an elliptic curve without complex multiplication and let $p$ be a prime such that $\operatorname{Im}\bar{\rho}_{E,p}$ is conjugate to a subgroup of $C_{ns}^{+}(p)$ and $\mathbb{Q}(E[p])/\mathbb{Q}$ is a nilpotent extension. Then $p$ is a Mersenne prime.
@@
In Table 2 we give models for all modular curves of prime level $p\in\{2,3,5,7\}$ for which $\mathbb{Q}(E[p])/\mathbb{Q}$ is nilpotent. These modular curves are isomorphic to $\mathbb{P}^{1}$ and hence there are infinitely many rational $j$-invariants of elliptic curves $E/\mathbb{Q}$ for which $\mathbb{Q}(E[p])/\mathbb{Q}$ is nilpotent.
@@ 4.5. The case of complex multiplication
One of the interesting properties of elliptic curves with complex multiplication is that their mod $p$ representations almost always have their images in the normalizer of a Cartan subgroup. Proposition 1.14(i) and (ii) of [45] state the following.
@@
[…]
The point of this proposition is that in these cases the mod $p$ images is as large as possible. This is useful because we know that in these cases the image of $\operatorname{Im}\bar{\rho}_{E,p}$ of in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is dihedral by Proposition 2.1 and has easily computable size. Remark 4.9 and Example 4.6 now imply that $\operatorname{Im}\bar{\rho}_{E,p}$ is nilpotent exactly when its image in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is a $2$-group.
@@
The discussion proceeding this theorem covers the case where $p\nmid{\rm disc}(\mathcal{O}).$ To handle the case where $p\mid{\rm disc}(\mathcal{O}),$ we refer to [45, Theorem 1.14] which shows that in this case $\operatorname{Im}\bar{\rho}_{E,p}$ isomorphic to one of the following groups
@@
In all three cases, the matrix $\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$ is in $\operatorname{Im}\bar{\rho}_{E,p}$ and so $\operatorname{Im}\bar{\rho}_{E,p}$ is not nilpotent by Proposition 4.10. ∎
@@ 4.5.1. The case when $E$ has complex multiplication by $\mathbb{Z}[\zeta_{3}]$
The images of the mod $p$ representations of $E$ depend on the value of $d$ modulo 6th powers. This relationship is explicitly classified in [45, Propositions 1.15 and 1.16]. We summarize the relevant parts of those propositions here for the convenience of the reader.
@@
[45, Propositions 1.15 and 1.16] Let $E/\mathbb{Q}$ be an elliptic curve with complex multiplication by $\mathbb{Z}[\zeta_{3}]$. Then the curve $E$ can be given by a Weierstrass equation of the form
@@
    If $d$ is a cube, then $\operatorname{Im}\bar{\rho}_{E,2}=\left\langle\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right\rangle$. Otherwise, $\operatorname{Im}\bar{\rho}_{E,2}=\operatorname{GL}_{2}(\mathbb{Z}/2\mathbb{Z})$.
@@
    If $4d$ is not a cube, then $\operatorname{Im}\bar{\rho}_{E,3}$ is conjugate to
[…]
    On the other hand, if $4d$ is a cube, then $\operatorname{Im}\bar{\rho}_{E,3}$ is conjugate to
[…]
@@
[…]
@@
    Suppose that $p\equiv 2$ or $5\bmod 9$ and let $e\in\{1,2\}$ such that $e\equiv\frac{p+1}{3}\bmod 3$. If $d\not\equiv 16p^{e}\bmod(\mathbb{Q}^{\times})^{3}$, then $\operatorname{Im}\bar{\rho}_{E,p}$ is conjugate to $C_{ns}^{+}(p)$. If $d\equiv 16p^{e}\bmod(\mathbb{Q}^{\times})^{3}$, then $\operatorname{Im}\bar{\rho}_{E,p}$ is conjugate in $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ to the subgroup generated by the unique index 3 subgroup of $C_{ns}(p)$ and $\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$.
The take away from this theorem is that for $E_{d}\colon y^{2}=x^{3}+d$, whether or not $\operatorname{Im}\bar{\rho}_{E,p}$ is nilpotent is completely controlled by $d\bmod(\mathbb{Q}^{\times})^{3}.$
For example, condition (1) tells us that $\mathbb{Q}(E_{d}[2])/\mathbb{Q}$ is a nilpotent extension exactly when $d$ is a cube. Similarly, condition (2) says that $\mathbb{Q}(E_{d}[3])/\mathbb{Q}$ is nilpotent when $4d$ is a cube.
We note that there are no Fermat primes $\equiv 1\pmod{9}$. By Remark 4.9 we have that $\operatorname{Im}\bar{\rho}_{E,p}$ is nilpotent if and only if its image in $\operatorname{PGL}_{2}(\mathbb{F}_{p})$, namely a dihedral group of order 2(p1)2(p-1), is nilpotent. By Example 4.6 this cannot occur since 2(p1)2(p-1) is not a power of $2$.
Thus if $p\equiv 1\pmod{9}$ the image of $\operatorname{Im}\bar{\rho}_{E,p}$ in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is dihedral if and only if
so condition (3) never yields a nilpotent $\mathbb{Q}(E[p])/\mathbb{Q}$. Likewise, there are no Mersenne primes $p\equiv 8\pmod{9}$.
The last cases that we have to deal with are the special cases that arise in cases (5) and (6) of Theorem 4.19. In cases (5) and (6) respectively, the image of $\bar{\rho}_{E,p}$ is contained in an index 3 subgroup of $C_{s}^{+}(p)$ and $C_{ns}^{+}(p)$ respectively. If we are in condition (5), then image of $\operatorname{Im}\bar{\rho}_{E,p}$ inside of $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is a dihedral group of size $\frac{2(p-1)}{3}$, while in condition (6) the image of $\operatorname{Im}\bar{\rho}_{E,p}$ inside of $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is a dihedral group of size $\frac{2(p+1)}{3}.$ This along with our previous analysis gives the following proposition.
@@
[…]
@@
Now we prove Corollary 1.10.
Proof.
If $E/\mathbb{Q}$ is an elliptic curve for which $\mathbb{Q}(E[19])/\mathbb{Q}$ is nilpotent then by Proposition 4.10, $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in the normalizer of a split or non-split Cartan subgroup. If $E$ is non-CM, the split Cartan case cannot occur by Proposition 4.12 and the non-split Cartan case cannot occur by Proposition 4.14. If $E$ has CM by an order $\mathcal{O}$, then Proposition 4.18 forces $\mathcal{O}=\mathbb{Z}[\zeta_{3}]$. However, by Proposition 4.20 only cases where $p=3\cdot 2^{k}\pm 1$ can occur and $19$ does not have this form.
Proposition 4.20 allows us to construct elliptic curves $E_{d}\colon y^{2}=x^{3}+d$ for which $\mathbb{Q}(E_{d}[p])/\mathbb{Q}$ is nilpotent for $p=2$, $p=3$, $p=5=3\cdot 2-1$, $p=7=3\cdot 2+1$, $p=11=3\cdot 2^{2}-1$ and $p=13=3\cdot 2^{2}+1$. The prime $p=17$ is a Fermat prime which splits in $\mathbb{Z}[i]$ and so if $E:y^{2}=x^{3}-x$, then $\mathbb{Q}(E[17])/\mathbb{Q}$ is nilpotent by Proposition 4.18. ∎